# Match, Compare, or Select? An Investigation of Large Language Models for Entity Matching

### v2

@@ Abstract
Entity matching (EM) is a critical step in entity resolution (ER). Recently, entity matching based on large language models (LLMs) has shown great promise. However, current LLM-based entity matching approaches typically follow a binary matching paradigm that ignores the global consistency between record relationships. In this paper, we investigate various methodologies for LLM-based entity matching that incorporate record interactions from different perspectives. Specifically, we comprehensively compare three representative strategies: matching, comparing, and selecting, and analyze their respective advantages and challenges in diverse scenarios. Based on our findings, we further design a compound entity matching framework (ComEM) that leverages the composition of multiple strategies and LLMs. ComEM benefits from the advantages of different sides and achieves improvements in both effectiveness and efficiency. Experimental results on 8 ER datasets and 9 LLMs verify the superiority of incorporating record interactions through the selecting strategy, as well as the further cost-effectiveness brought by ComEM.
@@
Tianshu Wang1,2, Xiaoyang Chen3, Hongyu Lin1, Xuanang Chen1, Xianpei Han1,4, Hao Wang5, Zhenyu Zeng5, Le Sun1,4 1Chinese Information Processing Laboratory, Institute of Software, Chinese Academy of Sciences 2Hangzhou Institute for Advanced Study, University of Chinese Academy of Sciences 3University of Chinese Academy of Sciences 4State Key Laboratory of Computer Science, Institute of Software, Chinese Academy of Sciences 5Alibaba Cloud Intelligence Group {tianshu2020, hongyu, xuanang2020, xianpei, sunle}@iscas.ac.cn, chenxiaoyang19@mails.ucas.ac.cn, cashenry@126.com, zhenyu.zzy@alibaba-inc.com
@@ 1 Introduction
Entity resolution (ER), also known as record linkage (Fellegi and Sunter, 1969) or deduplication (Elmagarmid et al., 2007), aims to identify and merge records that refer to the same real-world entity. Entity matching (EM) is a critical step in entity resolution that uses complex techniques to identify matching records from candidate pairs filtered by the blocking step (Papadakis et al., 2021). The recent emergence of large language models (LLMs) has introduced a new zero- or few-shot paradigm to EM, showing great promise (Narayan et al., 2022; Peeters and Bizer, 2023b, a; Fan et al., 2023; Li et al., 2024). LLM-based entity matching methods can achieve similar or even better performance than deep learning methods trained on large amounts of data, and are less susceptible to the unseen entity problem (Wang et al., 2022; Peeters et al., 2024).
@@
However, current LLM-based entity matching methods identify matches by classifying each pair of records independently. This matching strategy ignores the global consistency between record relationships and thus leads to suboptimal results. On the one hand, entity resolution requires more than independent classification due to the interconnected nature of record relationships (Getoor and Machanavajjhala, 2012). For example, in record linkage (i.e., clean-clean ER), a single record from one data source typically matches at most one record from another data source, since there are usually no duplicates in a single database. Unfortunately, matching-based approaches do not take advantage of this nature of record linkage. On the other hand, this strategy ignores the capabilities of LLMs to handle multiple records simultaneously to distinguish similar records. Using the records in Figure 1(c) as an example, if “Cruzer Glide”, “Cruzer Force”, and “Extreme Pro” appear in different records of the same context, LLMs are more likely to recognize that they are different SanDisk flash drive models, which helps with accurate matching. As a result, the matching strategy cannot fully unleash the potential of LLMs in EM.
@@
In this paper, we thoroughly investigate three strategies for LLM-based entity matching that incorporate record interactions from different perspectives, as shown in Figure 1. Specifically, apart from the conventional matching strategy shown in Figure 1(a), we investigate two additional strategies that leverage information from other records: 1) the comparing strategy, which identifies the record out of two that is more likely to match the anchor record, as shown in Figure 1(b); 2) the selecting strategy, which directly chooses the record from a list that is most likely to match the anchor record, as shown in Figure 1(c). Our research suggests that for LLM-based entity matching, incorporating record interactions is critical and can significantly improve EM performance in various scenarios. Therefore, the global selecting strategy is often the most effective. Nevertheless, we also observe that the selection accuracy decreases greatly as the position of the matching record increases in the candidate list. The position bias and limited long context understanding of current LLMs hinder the generality of the selecting strategy.
@@
Based on our findings, we design a compound entity matching framework (ComEM) that leverages the composition of multiple strategies and LLMs. Specifically, given an entity record and its $n$ potential matches obtained from the blocking step, we first preliminarily rank and filter these candidates using the local matching or comparing strategy, implemented with a medium-sized LLM. We then perform fine-grained identification on only the top $k$ candidates using the global selecting strategy, facilitated by a more powerful LLM. This approach not only mitigates the challenges and biases faced by the selecting strategy with too many options, but also reduces the cost of LLM invocations caused by composing multiple strategies. Consequently, by integrating the advantages of different strategies and LLMs, ComEM achieves a more effective and efficient entity matching process.
@@
To investigate different strategies and to evaluate our ComEM framework, we conducted in-depth experiments on 8 ER datasets. Experimental results verify the effectiveness of incorporating record interactions through the selecting strategy, with an average 17.58% improvement in F1 over the current matching strategy. In addition, we examined the effect of 9 different LLMs using these strategies on identification or ranking. Based on the results, ComEM is able to further improve the average F1 of the single selecting strategy by up to 4.01% while reducing the cost.
Contributions.
Generally speaking, our contributions can be summarized as follows [^1] The source code of this paper is available at: anonymous.4open.science/r/LLM4EM and supplementary material.:
@@ 2.1 Entity Resolution
As the core of data integration and cleaning, entity resolution has received extensive attention over the past decades (Fellegi and Sunter, 1969; Getoor and Machanavajjhala, 2012; Binette and Steorts, 2020; Papadakis et al., 2021). The blocking-and-matching pipeline has become the mainstream of entity resolution, where blocking filters out obviously dissimilar records and matching identifies duplicates through complex techniques.
Blocking.  Traditional blocking approaches group records into blocks by shared signatures, followed by cleaning up unnecessary blocks and comparisons (Papadakis et al., 2022). Meta-blocking further reduces superfluous candidates by weighting potential record pairs and graph pruning (Papadakis et al., 2014). Recently, nearest-neighbor search techniques, especially cardinality-based ones, have gained more attention and achieved state-of-the-art (SOTA) results for blocking (Thirumuruganathan et al., 2021; Paulsen et al., 2023).
Entity Matching.  The open and complex nature of entity matching has spurred the development of various approaches to address this persistent challenge, including rule-based (Benjelloun et al., 2009; Li et al., 2015), distance-based (Bilenko et al., 2003), and probabilistic methods (Fellegi and Sunter, 1969; Wu et al., 2020), etc. With the advent of deep learning methods (Mudgal et al., 2018), especially pre-trained language models (PLMs) (Li et al., 2020), entity matching has made significant progress (Barlaug and Gulla, 2021; Tu et al., 2023). The emergence of LLMs brings a new zero- or few-shot paradigm to entity matching (Narayan et al., 2022; Peeters and Bizer, 2023a), alleviating training data requirements. Most deep learning and LLM-based approaches treat entity matching as an independent classification problem, except for GNEM (Chen et al., 2021), which models this task as a collective classification task on graphs. To the best of our knowledge, this is the first effort to formulate entity matching as a comparison or selection task using LLMs.
@@ 2.2 Large Language Model
The advent of LLMs such as ChatGPT marks a significant advance in artificial intelligence, offering unprecedented natural language understanding and generation capabilities, and even general intelligence (Bubeck et al., 2023). By scaling up the model and data size of PLMs, LLMs exhibit emergent abilities (Wei et al., 2022) and can thus solve a variety of complex tasks by “prompt engineering” without “fine-tuning”. For more technical details on LLMs, we refer the reader to the related survey (Zhao et al., 2023).
@@ 3.1 Problem Formulation
We formulate the task of entity matching as the process of identifying matching records from a given entity record $r$ and its $n$ potential matches $R=\{r_{1},r_{2},\ldots,r_{n}\}$ obtained from blocking. This formulation mitigates the limitations of independent pairwise matching and fits real-world entity resolution scenarios. First, current SOTA blocking methods adhere to the k-nearest neighbor (kNN) search paradigm, which retrieves a list of potential matches for each entity record, rather than generating candidate matches pairwise as in traditional blocking workflows. In addition, this formulation accommodates both single-source deduplication and dual-source record linkage, and makes good use of the 1-1 assumption, i.e., record $r$ matches at most one of the potential matches $R$. This assumption is widespread in record linkage, and deduplication with canonical entity construction.
@@ 3.2 LLM as a Matcher
Recent work formulates entity matching as a binary classification task based on LLMs (Narayan et al., 2022; Peeters and Bizer, 2023b, a; Fan et al., 2023; Li et al., 2024). In this strategy, an LLM acts as a pairwise matcher to determine whether two records match. Specifically, given an entity record $r$ and its potential matches $R=\{r_{1},r_{2},\ldots,r_{n}\}$, this approach independently classifies each pair of records $(r,r_{i})_{1\leq i\leq n}$ as matching or not by interfacing LLMs with an appropriate matching prompt, as shown in Figure 1(a):
@@
Unlike previous studies, the core of LLM-based applications is to prompt LLMs to generate the correct answer, namely prompt engineering. An appropriate prompt should include the task instruction, such as “Do these two records refer to the same real-world entity? Answer Yes or No”. Optionally, a prompt could include detailed rules or several in-context learning examples to guide LLMs in performing this task. Given the need for long contexts in other strategies, and the instability of existing prompt engineering methods for entity matching Peeters and Bizer (2023a), we only attempt few-shot prompting for the matching strategy and leave the exploration of better prompt engineering with different strategies to future work.
This independent matching strategy ignores the global consistency of ER, as well as the capabilities of LLMs to incorporate record interactions. For record linkage, according to the well-known 1-1 assumption, each entity record $r$ matches at most one record of the potential matches $R$. For deduplication, this assumption can also be satisfied by constructing canonical entities. The traditional solution to satisfy these constraints is to construct a graph based on the similarity scores $s_{i}$ of record pairs $(r,r_{i})$ and to further cluster on the similarity graph. We can obtain the similarity scores from LLMs by calibrating the generated probabilities $p$ of labels Qin et al. (2023). Formally, the similarity score $s_{i}$ can be defined as:
@@ 3.3 LLM as a Comparator
where $\mathbb{1}_{r_{i}>_{r}r_{j}}$ and $\mathbb{1}_{r_{i}=_{r}r_{j}}$ indicate that record $r_{i}$ wins twice and once in comparison with record $r_{j}$ to record $r$. When LLMs do provide probabilities, the similarity scores $s_{i}$ can be defined as:
@@
3.4 LLM as a Selector
where $1,\ldots,n$ indicates the corresponding record.
In this way, LLMs can be explicitly required to identify only one match per record $r$ from the potential matches $R$. Furthermore, feeding LLMs all potential matches in the same context at a time allows LLMs to make better decisions by considering interactions between records. For example, if “SanDisk Cruzer Glide”, “SanDisk Cruzer Force”, and “SanDisk Extreme Pro” appear in different records of the same context, it is easier for LLMs to recognize that these are different model names of SanDisk flash drives and select the actual match.
One challenge in applying the selecting strategy to LLM-based entity matching is that there is often no actual match of record $r$ in potential matches $R$, because entity matching is an imbalanced task. A trivial solution to this challenge could be to perform a pairwise matching after the selection, which would undermine the advantages of the selecting strategy. Another method could be to add “none of the above” as an additional option to allow LLMs to refuse to select any record from the potential matches, which can be formulated as:
where 00 indicates the “none of the above” option.
However, the selecting strategy relies heavily on the capabilities of LLMs for fine-grained understanding and implicit ranking in long contexts. Our experimental results show that the current LLMs suffer from position bias, with the selection accuracy decreasing significantly as the position of the matching record increases in the candidate list (§ 4.3). In practice, the recall-oriented blocking step often generates a considerable number of potential matches for each record, exceeding the context length that LLMs can effectively reasoning (Levy et al., 2024). Therefore, it is a challenge to mitigate the position bias and the long context requirement for the selecting strategy.
3.5 Compound Entity Matching Framework
@@
Table 1: Comparison of different strategies. “” means that the matching strategy can only calibrate similarity scores if the generation probability is available. “# Input Records” represents the number of (#) records input to LLMs using different strategies for record $r$ and its $n$ potential matches $R$. $k$ denotes the number of top candidates considered by the comparing strategy.

### v3

@@ Abstract
Entity matching (EM) is a critical step in entity resolution (ER). Recently, entity matching based on large language models (LLMs) has shown great promise. However, current LLM-based entity matching approaches typically follow a binary matching paradigm that ignores the global consistency among record relationships. In this paper, we investigate various methodologies for LLM-based entity matching that incorporate record interactions from different perspectives. Specifically, we comprehensively compare three representative strategies: matching, comparing, and selecting, and analyze their respective advantages and challenges in diverse scenarios. Based on our findings, we further design a compound entity matching framework (ComEM) that leverages the composition of multiple strategies and LLMs. ComEM benefits from the advantages of different sides and achieves improvements in both effectiveness and efficiency. Experimental results on 8 ER datasets and 10 LLMs verify the superiority of incorporating record interactions through the selecting strategy, as well as the further cost-effectiveness brought by ComEM.
@@
Tianshu Wang1,2, Xiaoyang Chen3, Hongyu Lin1, Xuanang Chen1thanks: Corresponding author., Xianpei Han1,4footnotemark: , Hao Wang5, Zhenyu Zeng5, Le Sun1,4 1Chinese Information Processing Laboratory, Institute of Software, Chinese Academy of Sciences 2Hangzhou Institute for Advanced Study, University of Chinese Academy of Sciences 3University of Chinese Academy of Sciences 4State Key Laboratory of Computer Science, Institute of Software, Chinese Academy of Sciences 5Alibaba Cloud Intelligence Group {tianshu2020, hongyu, chenxuanang, xianpei, sunle}@iscas.ac.cn, chenxiaoyang19@mails.ucas.ac.cn, cashenry@126.com, zhenyu.zzy@alibaba-inc.com
@@ 1 Introduction
Entity resolution (ER), also known as record linkage (Fellegi and Sunter, 1969) or deduplication (Elmagarmid et al., 2007), aims to identify and canonicalize records that refer to the same real-world entity. ER is a fundamental task of data integration and cleansing, with broad applications in maintaining data consistency, accurate data analysis, and informed decision making. Entity matching (EM) serves as a critical step in entity resolution that uses complex techniques to identify matching records from potential matches filtered by the blocking step (Papadakis et al., 2021). The recent emergence of large language models (LLMs) has introduced a new zero- or few-shot paradigm to EM, showing great promise (Narayan et al., 2022; Peeters and Bizer, 2023; Fan et al., 2024; Li et al., 2024; Peeters et al., 2025). LLM-based entity matching methods can achieve similar or even better performance than deep learning methods trained on large amounts of data, and are less susceptible to the unseen entity problem (Wang et al., 2022; Peeters et al., 2024).
@@
However, current LLM-based entity matching methods identify matches by classifying each pair of records independently. This matching strategy ignores the global consistency [^1] We refer to the interdependence of matching decisions in ER as global consistency. See Appendix A for more details. among record relationships and thus leads to suboptimal results. On the one hand, entity resolution requires more than independent classification due to the interconnected nature of record relationships (Getoor and Machanavajjhala, 2012). For example, in record linkage (i.e., clean-clean ER), a single record from one data source typically matches at most one record from another data source, since there are usually no duplicates in a single database. Unfortunately, matching-based approaches do not take advantage of this nature of record linkage. On the other hand, this strategy ignores the capabilities of LLMs to handle multiple records simultaneously to distinguish similar records. Using the records in Figure 1(c) as an example, if “Cruzer Glide”, “Cruzer Force”, and “Extreme Pro” appear in different records of the same context, LLMs are more likely to recognize that they are different SanDisk flash drive models, which helps with accurate matching. As a result, the matching strategy cannot fully unleash the potential of LLMs in EM.
@@
In this paper, we thoroughly investigate three strategies for LLM-based entity matching that incorporate record interactions from different perspectives, as shown in Figure 1. Specifically, apart from the conventional matching strategy shown in Figure 1(a), we investigate two additional strategies that leverage information from other records: 1) the comparing strategy, which identifies the record out of two that is more likely to match the anchor record, as shown in Figure 1(b); 2) the selecting strategy, which directly chooses the record from a list that is most likely to match the anchor record, as shown in Figure 1(c). Our research suggests that for LLM-based entity matching, incorporating record interactions is critical and can significantly improve entity matching performance in various scenarios. Among these strategies, the selecting strategy is often the most cost-effective. Nevertheless, we also observe that the selection accuracy varies significantly as the position of the matching record increases in the candidate list. The position bias and limited long context understanding of current LLMs Levy et al. (2024) hinder the generality of the selecting strategy.
@@
Based on our findings, we design a compound entity matching framework (ComEM) that leverages the composition of multiple strategies and LLMs. Specifically, given an entity record and its $n$ potential matches obtained from the blocking step, we first preliminarily rank and filter these candidates using the local matching or comparing strategy, implemented with a medium-sized LLM. We then perform fine-grained identification on only the top $k$ candidates using the global selecting strategy, facilitated by a more powerful LLM. This approach not only mitigates the challenges and biases faced by the selecting strategy with too many options, but also reduces the cost of LLM invocations caused by composing multiple strategies. Consequently, by integrating the advantages of different strategies and LLMs, ComEM achieves a more effective and cost-efficient entity matching process.
@@
To investigate different strategies and to evaluate ComEM, we conducted in-depth experiments on eight ER datasets. Experimental results verify the effectiveness of incorporating record interactions through the selecting strategy, with an average 16.02% improvement in F1 over the current matching strategy. In addition, we examined the effect of 10 different LLMs using these strategies on identification or ranking. Ultimately, ComEM is able to further improve average F1 of the single selecting strategy by up to 4% while reducing the cost.
Contributions.  In general, our contributions can be summarized as follows [^2] Our code is available at github.com/tshu-w/ComEM to facilitate reproduction of our results.:
@@ 2 Related Work
Records Title Authors Venue Year
Anchor Lineage Tracing for General Data Warehouse Transformations Yingwei Cui, Jennifer Widom VLDB 2001
Potential Matches Lineage Tracing for General Data Warehouse Transformations Yingwei Cui, Jennifer Widom VLDB Journal 2003
Tracing the lineage of view data in a warehousing environment Yingwei Cui, Jennifer Widom, et al. TODS 2000
Lineage tracing for general data warehouse transformations Y. Cui, J. Widom VLDB 2001
Table 1: Example of our formulation for entity matching: Given an anchor record, identify the matching record (if any) from its potential matches. This example is taken from the DBLP-ACM dataset.
@@ 2.1 Entity Resolution
Entity resolution has received extensive attention over the past decades (Fellegi and Sunter, 1969; Getoor and Machanavajjhala, 2012; Papadakis et al., 2021; Binette and Steorts, 2022). The blocking-and-matching pipeline has become the mainstream of entity resolution, where blocking filters out obviously dissimilar records and matching identifies duplicates through complex techniques.
Blocking.  Traditional blocking approaches group records into blocks by shared signatures, followed by cleaning up unnecessary blocks and comparisons (Papadakis et al., 2022). Meta-blocking further reduces superfluous candidates by weighting potential record pairs and graph pruning (Papadakis et al., 2014). Recently, nearest-neighbor search techniques, especially cardinality-based ones, have gained more attention and achieved state-of-the-art (SOTA) results (Thirumuruganathan et al., 2021; Paulsen et al., 2023; Wang and Zhang, 2024).
Entity Matching.  The open and complex nature of entity matching has spurred the development of various approaches to address this persistent challenge, including rule-based (Benjelloun et al., 2009; Li et al., 2015), distance-based (Bilenko et al., 2003), and probabilistic methods (Fellegi and Sunter, 1969; Wu et al., 2020), etc. With the advent of deep learning methods (Mudgal et al., 2018), especially pre-trained language models (PLMs) (Li et al., 2020), entity matching has made significant progress (Barlaug and Gulla, 2021; Tu et al., 2023; Wu et al., 2023; Zhang et al., 2024). The emergence of LLMs brings a new zero- or few-shot paradigm to entity matching (Narayan et al., 2022; Xu et al., 2024; Peeters et al., 2025), alleviating training data requirements. Most deep learning and LLM-based approaches treat entity matching as an independent binary classification problem, except for GNEM (Chen et al., 2021), which models this task as a collective classification task on graphs.
@@ 2.2 Large Language Model
The advent of LLMs such as ChatGPT marks a significant advance in artificial intelligence, offering unprecedented natural language understanding and generation capabilities. By scaling up the model and data size of PLMs, LLMs exhibit emergent abilities (Wei et al., 2022) and can thus solve a variety of complex tasks by prompt engineering. For more technical details on LLMs, we refer the reader to the related survey (Zhao et al., 2023).
While LLMs have shown promising results in classification and ranking tasks (Sun et al., 2023; Qin et al., 2024), applying LLMs to entity matching presents unique challenges and opportunities. Our work differs from previous research in three aspects: First, we propose a novel paradigm that formulates entity matching as a comparison or selection task. Second, we demonstrate that the effectiveness of pairwise and listwise strategies in entity matching exhibits different patterns compared to ranking. Finally, through a comprehensive cross-model and cross-strategy evaluation, we reveal several key insights about LLM-based EM, which motivate the design of our ComEM framework.
@@ 3.1 Problem Formulation
We formulate the task of entity matching as the process of identifying matching records from a given anchor record $r$ and its $n$ potential matches $R=\{r_{1},r_{2},\ldots,r_{n}\}$ obtained from blocking, as illustrated in Table 1. This formulation mitigates the limitations of independent pairwise matching and fits real-world ER scenarios. First, current SOTA blocking methods adhere to the k-nearest neighbor (kNN) search paradigm, which retrieves a list of potential matches for each entity record. In addition, this formulation accommodates both single-source deduplication and dual-source record linkage, and makes good use of the one-to-one assumption, i.e., record $r$ matches at most one of the records in potential matches $R$. This assumption is widespread in record linkage, and deduplication with canonicalization.
@@ 3.2 LLM as a Matcher
Recent work formulates entity matching as a binary classification task based on LLMs (Narayan et al., 2022; Peeters and Bizer, 2023; Fan et al., 2024; Li et al., 2024; Peeters et al., 2025). In this strategy, an LLM acts as a pairwise matcher to determine whether two records match. Specifically, given an entity record $r$ and its potential matches $R=\{r_{1},r_{2},\ldots,r_{n}\}$, this approach independently classifies each pair of records $(r,r_{i})_{1\leq i\leq n}$ as matching or not by interfacing LLMs with an appropriate matching prompt, as shown in Figure 1(a):
@@
Unlike previous studies, the core of LLM-based applications is to prompt LLMs to generate the correct answer, namely prompt engineering. An appropriate prompt should include the task instruction, such as “Do these two records refer to the same real-world entity? Answer Yes or No”. Optionally, a prompt could include detailed rules or several in-context learning examples to guide LLMs in performing this task. Given the need for long contexts in other strategies, and the instability of existing prompt engineering methods for entity matching Peeters et al. (2025), we only attempt few-shot prompting for the matching strategy and leave the exploration of better prompt engineering with different strategies to future work.
This independent matching strategy ignores the global consistency of ER, as well as the capabilities of LLMs to incorporate record interactions. The traditional solution to satisfy these constraints is to construct a graph based on the similarity scores $s_{i}$ of record pairs $(r,r_{i})$ and to further cluster on the similarity graph. We can obtain the similarity scores from LLMs by calibrating the generated probabilities $p$ of labels Liang et al. (2023). Formally, the similarity score $s_{i}$ can be defined as:
@@ 3.3 LLM as a Comparator
where $\mathbb{1}_{r_{i}>_{r}r_{j}}$ and $\mathbb{1}_{r_{i}=_{r}r_{j}}$ indicate that record $r_{i}$ wins twice and once in comparison with record $r_{j}$ to record $r$. When LLMs do provide probabilities, the similarity score $s_{i}$ can be defined as:
@@
Table 2: Comparison of different strategies. “” means that the matching strategy can only calibrate similarity scores if the generation probability is available. “# LLM Invocations” and “# Input Records” represent the number of (#) LLM invocations and records input to LLMs using different strategies for record $r$ and its $n$ potential matches $R$, respectively. $k$ denotes the number of top candidates considered by the comparing strategy.
3.4 LLM as a Selector
where $1,\ldots,n$ indicates the corresponding record.
In this way, LLMs can be explicitly required to identify only one match per record $r$ from the potential matches $R$. Furthermore, feeding LLMs all potential matches in the same context at a time allows LLMs to make better decisions by considering interactions between records. Using Table 1 as an example, it is easier for LLMs to recognize the less critical attributes, such as authors, and identify the third record as the true match by comparing the values of title and year across different records.
One challenge in applying the selecting strategy to LLM-based entity matching is that sometimes there is no actual match of record $r$ in potential matches $R$. A trivial solution to this challenge could be to perform a pairwise matching after the selection, which would undermine the advantages of the selecting strategy. Another method could be to add “none of the above” as an additional option to allow LLMs to refuse to select any record from the potential matches, which can be formulated as:
where 00 indicates the “none of the above” option.
However, the selecting strategy relies heavily on the capabilities of LLMs for fine-grained understanding and implicit ranking in long contexts. Our experimental results show that current LLMs suffer from position bias, with the selection accuracy varying significantly as the position of the matching record increases in the candidate list (§ 4.3). In practice, the recall-oriented blocking step often generates a considerable number of potential matches for each record, exceeding the context length that LLMs can effectively reason (Levy et al., 2024). Therefore, it is a challenge to mitigate the position bias and the long context requirement for the selecting strategy.
3.5 Compound Entity Matching Framework
Figure 2: Illustration of ComEM. It first filters candidate records by matching or comparing strategies and then identifies the match via the selecting strategy.